NASA going to launch a rocket to Mars, with the rocket blasting off from the Earth when the Earth and Mars are aligned along a straight line from the Sun. If the Mars is now 60° ahead of the Earth in its orbit around the sun, when should NASA launch the rocket? All the planets orbit the sun in the same direction and 1 year on Mars is 1.9 Earth-years. Assume circular orbits for the Earth and the Mars.


1
Expert's answer
2021-06-21T15:07:21-0400

Apply"\theta=wt"

For alignment the earth must move through 60° more than mars in the same time t

"w_c=\frac{360°}{yr}"

"w_m=\frac{360°}{1.9yr}"

"\theta_c=\theta_m+60°"

"w_ct=w_mt+60\u00b0"

"w_c=\frac{60°}{w_c-w_m}"

"t=\frac{60°}{\frac{360}{1yr}-\frac{360°}{1.9yr}}"

"t=\frac{60°}{360}(\frac{1}{\frac{0.9yr}{1.9yr^2}})"

"t=0.352yr=128days"

Earth has a larger angular velocity then mass and completes one orbit in less time